Thermal Rising, Inc., makes paragliders for sale through specialty sporting goods stores. The company has a standard paraglider model, but also makes custom-designed paragliders. Management has designed an activity-based costing system with the following activity cost pools and activity rates:

Activity Cost Pool Activity Rate  

Supporting direct labor $ 20 per direct labor-hour

Order processing    $ 198 per order

Custom design processing $ 254 per custom design

Customer service $ 418 per customer

Management would like an analysis of the profitability of a particular customer, Big Sky Outfitters, which has ordered the following products over the last 12 months:

Standard Model Custom Design

Number of gliders 14 2

Number of orders 2   2

Number of custom designs 0 2

Direct labor-hours per glider 27.50 33.00

Selling price per glider $ 1,650 $ 2,360

Direct materials cost per glider $ 454 $ 582

The company’s direct labor rate is $20 per hour.

Required: Using the company’s activity-based costing system, compute the customer margin of Big Sky Outfitters. (Round your intermediate calculations and final answer to the nearest whole dollar amount. Loss amounts should be entered with a minus sign.)
Answer #1
Revenue 27820 =(1650*14)+(2360*2)
Costs:
Direct materials 7520 =(454*14)+(582*2)
Direct labor 9020 =(27.5*20*14)+(33*20*2)
Supporting Direct labor 9020 =(27.5*20*14)+(33*20*2)
Order processing 792 =(2*198)+(2*198)
Custom design processing 508 =(2*254)
Customer service 418
Total costs 27278
Customer margin 542